DIRC Operations Weekly Report

Period: Wednesday 2007/6/6 - Wednesday 2007/6/13


Oncall: Brad Wogsland
Fast Monitoring: Jose Benitez, Justine Serrano, Brad Wogsland

Contents


    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~1 x 1034 (last week: ~9.5 x 1033)

Run 6 PEP delivery: 3.49 fb-1 (last weeks: 4.8 fb-1)

BaBar integrated lumi:3.42 fb-1 including 1.55 fb-1 Offpeak (last weeks: 4.6 fb-1)

Run 6 total delivery: 46.15 fb-1

Run 6 total BaBar recorded: 44.43 fb-1 including 1.55 fb-1 off-peak data

PEP activities/issues this week

  • Generally delivered decent luminosity but with high backgrounds.
  • Switched Monday to Offpeak running.
  • Hot weather is starting to disrupt datatime operations.

    BaBar activities/issues this week

  • UPS battery replacement.

    DIRC activities/issues this week

  • Paged due for background in the DIRC on 6/8 and 6/9.
  • New problem with 8 flaky channels in sector 3.
  • Masked 8 noisy channels (7 in sector 6 and 1 in sector 8)